# Hierarchical Context Alignment with Disentangled Geometric and Temporal Modeling for Semantic Occupancy Prediction

### v1

@@ Abstract
Index Terms:
3D visual perception, semantic occupancy prediction, hierarchical context alignment.
1 Introduction
@@
Comprehending holistic 3D scenes is crucial for autonomous driving systems, as it significantly influences the planning and obstacle avoidance capabilities for autonomous vehicle safety and efficiency [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. However, the limitations of real-world sensors, including restricted fields of view and measurement noise, present substantial challenges. To overcome these difficulties, 3D semantic occupancy prediction (SOP) has been developed to simultaneously infer the geometry and semantics of the scenario from partial observations [12, 1, 13, 14, 15, 16, 5, 6].
@@
Given the inherent 3D nature, numerous semantic occupancy prediction (SOP) solutions rely on LiDAR for accurate location measurements [12, 1, 13, 14, 15]. Although LiDAR provides precise depth information, it inevitably introduces significant cost and manual effort with dense annotations and sophisticated devices. Consequently, it is urgent to explore an efficient approach for precise SOP with a cost-effective scheme. This motivation has prompted the exploration of camera-based solutions, which are characterized by superior deployment efficiency and offer richer visual context, making them a promising alternative for SOP [16, 7, 17, 6].
@@
To construct accurate occupancy representations, previous camera-based SOP methods have explored contextual feature aggregation from both geometric and temporal perspectives [18, 19, 20, 16, 21]. As shown in Figure 1 (a), prior geometric modeling approach (e.g., OccFormer [17]) typically employs geometric lifting in the voxel feature construction process for image-to-3D transformation. While the temporal modeling approach (e.g., VoxFormer-T [5]) utilizes temporal coherence by stacking multiple historical frames as supplements to the current stereo frame, as illustrated in Figure 1 (b). Despite these significant contributions, these methods failed to simultaneously address both geometric and temporal aspects and also trivially fused contextual information in a black-box manner [19, 18].
As a result, the existing SOP solutions inevitably face the misalignment issue during the scene modeling process. That is, the corresponding features at the same position across different frames may have different semantic meanings during the aggregation process, which can result in fuzzy contextual fusion and unstable representation learning in camera-based visual perception [22, 23, 24]. As illustrated in Figure 2, such a misalignment issue could lead to unreliable prediction results and unstable learning processes for semantic occupancy prediction. Specifically, the geometric modeling presented in OccFormer [17] neglects the uncertainty inherent in monocular depth estimation during the voxel feature lifting process. This oversight causes geometric ambiguity when depth information is integrated with the corresponding contextual features. Besides, the temporal modeling in VoxFormer-T [5] often assumes that temporal features from different viewpoints directly correspond at the pixel level through simple straightforward aggregation. This assumption ignores the positional changes of shared semantic content across different perspectives, leading to blurred predictive information and compromising the stability of SOP.
Figure 1: Our hierarchical context alignment learning method versus previous geometric modeling (e.g., OccFormer [17]) and temporal modeling (e.g., VoxFormer-T [5]) methods for semantic occupancy prediction.
@@
Figure 2: The effect of the hierarchical context alignment on the SemanticKITTI validation set. We remove both the temporal alignment and the geometric alignment to implement the setting of ’w/o align’. The proposed hierarchical context alignment strategy captures more reliable and comprehensive semantic scenes, and leads to more stable representation modeling in the learning process.
@@
To this end, we propose a novel Hierarchical context alignment SOP scheme, termed Hi-SOP, which disentangles the complex task of semantic scene comprehension into distinct geometric and temporal context modeling processes. These two branches are globally aligned and integrated to achieve reliable camera-based semantic occupancy prediction. As shown in Figure 1 (c), our method takes advantage of the complementary merits of the geometric and temporal representations (one for the spatial and the other for historic feature perception), and hierarchically aligns them for a reliable context composition. This hierarchical context alignment with a disentangled local-global structure is composed of two sequential steps: (I) individual geometric alignment across frames at different views with optional monocular or stereo depth estimation, and temporal alignment through homography warping and confidence-aware dynamic refinement; (II) globally align and compose the geometric and temporal context within a unified space through semantically consistent transformation and aggregation for a final reliable occupancy prediction. As shown in Figure 2, our proposed hierarchical context alignment strategy Hi-SOP shows promising performance in capturing more reliable and comprehensive semantic scenes, leading to more accurate perdition results and more stable learning processes.
@@
More specifically, to facilitate reliable geometric alignment, we design a Geometric Confidence-aware Lifting (GCL) module in the geometric alignment branch, which models geometric information with depth distribution confidence awareness before integrating it with corresponding contextual features for voxel feature lifting. For temporal-wise alignment, we first employ an epipolar homography-warping to explicitly align temporal invariant features and create temporal feature volumes to preserve detailed context. To separate critical relevant context from redundant information, we construct a Cross-frame Pattern Affinity (CPA) to measure the contextual relevance, and accordingly based on it refine the dynamic temporal content to compensate for incomplete observations. Finally, in the composition stage, we propose to globally align the geometric context with the temporal context within a unified space by a Depth-Hypothesis-Based Transformation (DHBT) for semantic-consistent aggregation, which takes the depth hypothesis of the temporal feature volume as the distance axis and employ voxel-pooling operation to splat the volumetric features into the unified space.
@@
We conduct extensive experiments to validate the advantages of our proposed hierarchical context alignment paradigm for Semantic Occupancy Prediction (SOP) in the semantic scene completion (SSC) and LiDAR semantic segmentation tasks. For SSC, our camera-based Hi-SOP outperforms state-of-the-art VoxFormer-T [5] on the SemanticKITTI [25] benchmark and even surpasses LiDAR-based methods on the NuScenes-Occupancy [26] benchmark. We also evaluate our method on the NuScenes [27] dataset for LiDAR semantic segmentation. Hi-SOP surpasses TPVFormer [6] with a relative improvement of 24.28% in terms of mIoU. This work is an extension version that upgrades our previous ECCV-24 conference paper HTCL [28] into Hi-SOP with the following new contributions:
@@
1) Conceptually, we introduce a new hierarchical context alignment paradigm that first disentangles geometric and temporal context learning into a local-global hierarchy, and then aggregates them based on semantics consistency for a complementary composition.
@@
3) Experimentally, the initial HTCL focuses only on the semantic scene completion task, while Hi-SOP extends the framework for semantic occupancy prediction tasks including semantic scene completion and LiDAR semantic segmentation. Our code and demo video are available in the supplementary material, and the project website is also available at https://arlo0o.github.io/hisop.github.io/.
2 Related Work
2.1 Semantic Occupancy Prediction
@@
Semantic Occupancy Prediction (SOP), also referred to as Semantic Scene Completion (SSC), represents a comprehensive 3D perception task that concurrently tackles semantic segmentation and scene completion [25, 29, 30, 31, 28]. Many studies have traditionally utilized LiDAR as the primary data source to capitalize on its 3D geometric information [12, 1, 13]. However, due to its cost-effectiveness and portability, camera-based 3D SOP has recently attracted significant interest [25, 32, 20, 33, 29, 1, 34, 35, 15, 14, 36, 37, 16, 5, 6]. MonoScene [16] pioneered the inference of geometry and semantics from a single RGB image using 2D-3D feature projection. Following this innovation, numerous studies have expanded the scope of camera-based 3D scene perception [6, 17, 7, 31]. OccFormer [17] employs a monocular depth net and context net to lift voxel feature volume, which is processed with a dual-path transformer block for semantic occupancy prediction. TPVFormer [6] introduces a tri-perspective view to enhance the detailed representation of a 3D scene. SurroundOcc [7] estimates dense 3D occupancy using multi-view image inputs. VPD [38] utilizes conditional diffusion models for 3D perception tasks, including multi-view stereo and semantic occupancy prediction. Nonetheless, these methods attempt to model complex 3D scenes using single-timestep images, which proves suboptimal for this inherently challenging problem due to the lack of comprehensive visual cues. In this paper, we advocate for leveraging reliable temporal data to dynamically integrate semantic context and compensate for incomplete observations.
2.2 Geometry Learning in BEV Representation
@@
The bird’s-eye view (BEV) is a prevalent representation in 3D object detection, offering a comprehensive depiction of layouts and strong hallucination capabilities from a top-down perspective [30, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48]. The Lift-Splat approach [30] initially introduced the extraction of BEV representations from multiple cameras by implicitly unprojecting 2D visual inputs through estimated depth distributions. To improve geometric modeling in the lifting process, BEVDepth [37] employs a camera-aware monocular depth estimation module, enhancing depth accuracy in BEV-based 3D detection. As an effective representation for 3D scenarios, BEV representations have also been effectively utilized in recent occupancy-based perception works [17, 7, 31]. Notably, StereoScene [31] takes advantage of stereo matching technology and utilizes stereo images to improve the geometric information in the BEV representation and achieves remarkable enhancements. In this study, we develop a new framework that incorporates monocular or stereo depth estimation to explicitly model geometric information with depth distribution confidence awareness. Moreover, the geometric context is aligned with the temporal context into a unified space through depth-hypothesis-based transformation for stable representation aggregation.
2.3 Temporal Modeling in 3D Visual Perception
@@
The incorporation of temporal information has gained prominence in applications such as temporal 3D object detection [49, 50, 51, 52, 53, 54, 55, 56] and video depth estimation [57, 58, 59, 60, 61, 62, 63], enhancing overall prediction accuracy. Temporal 3D object detection generally targets coarse-grained, regional-level predictions [56, 54], whereas video depth estimation techniques strive to establish correspondences across sequential video frames [58, 59]. Nevertheless, such approaches fall short in SOP, where capturing fine-grained features is crucial for dense semantic perception. VoxFormer-T [5] establishes the first temporal framework for camera-based SOP by merely stacking features from different frames, yet the temporal correspondence modeling for the dense perception task of SOP remains unexplored. In this paper, we propose to explicitly model the temporal context correlation through pattern affinity, thereby aggregating reliable aligned temporal content and mitigating the impact of incomplete observations.
@@
3 Methodology
3.1 Overview
3.1.1 Preliminary
Given a sequence of temporal RGB images $I_{{set}}^{{rgb}}=\{I_{t}^{{rgb}},I_{t-1}^{{rgb}},\ldots\}$, our objective is to estimate the semantic voxel grid for semantic scene completion or LiDAR segmentation [16, 17]. We focus on current and historical image frames, excluding future frames [5] to devise a practical method for real-world applications. The scene is represented as a voxel grid $\mathbf{V}$ with dimensions $\mathbb{R}^{H\times W\times Z}$, where $H$, $W$, and $Z$ denote the height, width, and depth of the grid, respectively. Each voxel within this grid is classified into one of the semantic categories in the set $\{c_{0},c_{1},\ldots,c_{N}\}$, where $c_{0}$ indicates empty space and $\{c_{1},c_{2},\ldots,c_{N}\}$ correspond to $N$ distinct semantic classes. With the proposed framework $\Theta$, we seek to learn a transformation defined as:

### v2

@@ Abstract
I Introduction
@@
Comprehending holistic 3D scenes is crucial for autonomous driving systems, as it significantly influences the planning and obstacle avoidance capabilities for autonomous vehicle safety and efficiency [li2023front, li2023voxformer, huang2023tri, wei2023surroundocc, li2023delving, luo2024exploring]. However, the limitations of real-world sensors, including restricted fields of view and measurement noise, present substantial challenges. To overcome these difficulties, 3D semantic occupancy prediction (SOP) has been developed to simultaneously infer the geometry and semantics of the scenario from partial observations [rist2021semantic, roldao2020lmscnet, cao2022monoscene, li2023voxformer, huang2023tri].
@@
Given the inherent 3D nature, numerous semantic occupancy prediction (SOP) solutions rely on LiDAR for accurate location measurements [rist2021semantic, garbade2019two, roldao2020lmscnet]. Although LiDAR provides precise depth information, it inevitably introduces significant cost and manual effort with dense annotations and sophisticated devices. Consequently, it is urgent to explore an efficient approach for precise SOP with a cost-effective scheme. This motivation has prompted the exploration of camera-based solutions, which are characterized by superior deployment efficiency and offer richer visual context, making them a promising alternative for SOP [cao2022monoscene, wei2023surroundocc, zhang2023occformer, huang2023tri].
@@
To construct accurate occupancy representations, previous camera-based SOP methods have explored contextual feature aggregation from both geometric and temporal perspectives [li2024bridging, xue2024bi, song2017semantic, cao2022monoscene, roldao20223d]. As shown in Figure 1 (a), prior geometric modeling approach (eg., OccFormer [zhang2023occformer]) typically employs geometric lifting in the voxel feature construction process for image-to-3D transformation. While the temporal modeling approach (eg., VoxFormer-T [li2023voxformer]) utilizes temporal coherence by stacking multiple historical frames as supplements to the current stereo frame, as illustrated in Figure 1 (b). Despite these significant contributions, these methods failed to simultaneously address both geometric and temporal aspects and also trivially fused contextual information in a black-box manner [xue2024bi, li2024bridging].
Figure 1: Comparison of our hierarchical context alignment method with prior geometric modeling (e.g., OccFormer [zhang2023occformer]) and temporal modeling approaches (e.g., VoxFormer-T [li2023voxformer]) for semantic occupancy prediction. Previous methods handle either geometric lifting or temporal stacking separately, often fusing contexts trivially in a black-box manner, which leads to misalignment and unreliable fusion. In contrast, our approach integrates both geometric and temporal representations in a hierarchically aligned manner, enabling robust contextual composition.
@@
Figure 2: (a) Visualization examples of the misalignment issue. The same car appears at different locations across multiple frames. Directly fusing contextual information from these frames neglects positional shifts and scale variations, which could lead to ambiguous contextual aggregation and unstable representation learning for semantic occupancy prediction. (b) The effect of the hierarchical context alignment on the SemanticKITTI validation set. We remove both the temporal alignment and the geometric alignment to implement the setting of ‘w/o align’. The proposed hierarchical context alignment strategy captures more reliable and comprehensive semantic scenes, and leads to more stable representation modeling in the learning process.
As a result, existing SOP solutions inevitably face the misalignment issue during the scene modeling process in the geometric or temporal aspects. Specifically, features corresponding to the same spatial location across different frames may carry inconsistent semantic meanings during aggregation, which can result in fuzzy contextual fusion and unstable representation learning in camera-based visual perception [song2025traf, pan2025overcoming, yang2024align, park2024adversarial, guo2024damsdet]. For instance, as illustrated in Figure 2, a moving car may appear at different locations in multiple RGB frames. Directly fusing feature representations extracted from these frames assumes pixel-level correspondence between features from different viewpoints, which is often invalid. This mismatch causes blurred predictive information, as shared semantic content undergoes unpredictable positional shifts that are not consistently aligned across perspectives. Notably, previous SOP methods [zhang2023occformer, li2024hierarchical2, li2023voxformer] fail to jointly address geometric and temporal aspects while naively fusing contextual information in a black-box manner, which exacerbates misalignment issues, resulting in unreliable contextual fusion and unstable representation learning. For instance, as shown in Figure 1 (a), the geometric modeling presented in OccFormer [zhang2023occformer] neglects the uncertainty inherent in monocular depth estimation during the voxel feature lifting process. This oversight causes geometric ambiguity when depth information is integrated with the corresponding contextual features. Similarly, Figure 1(b) highlights that VoxFormer-T [li2023voxformer] assumes direct pixel-level correspondence between temporal features from different viewpoints through simple aggregation. However, this assumption disregards positional shifts of shared semantic content across perspectives, yielding blurred predictions and compromising the stability of SOP.
@@
To address these limitations, we propose Hi-SOP, a novel hierarchical context alignment framework that addresses the critical misalignment issue in camera-based semantic occupancy prediction by disentangling and separately aligning geometric and temporal contexts before composition. These aligned contexts are then composed within a unified space to produce reliable semantic occupancy grids. As depicted in Figure 1 (c), our method takes advantage of the complementary merits of the geometric and temporal representations (one for the spatial and the other for historic feature perception), and hierarchically aligns them for a reliable context composition. This hierarchical context alignment with a disentangled local-global structure is composed of two sequential steps: (I) individual geometric alignment across frames at different views with optional monocular or stereo depth estimation, and temporal alignment through homography warping and affinity-aware dynamic refinement; (II) globally align and compose the geometric and temporal context within a unified space through semantically consistent transformation and aggregation for a final reliable occupancy prediction. As shown in Figure 2, our proposed hierarchical context alignment strategy Hi-SOP shows promising performance in capturing more reliable and comprehensive semantic scenes, leading to more accurate prediction results and more stable learning processes.
@@
More specifically, to facilitate reliable geometric alignment, we design a Geometric Confidence-aware Lifting (GCL) module in the geometric alignment branch, which models geometric information with depth distribution confidence awareness before integrating it with corresponding contextual features for voxel feature lifting. For temporal-wise alignment, we first employ an epipolar homography-warping to explicitly align temporal invariant features and create temporal feature volumes to preserve detailed context. To separate critical relevant context from redundant information, we construct a Cross-frame Pattern Affinity (CPA) to measure the contextual relevance, and accordingly based on it refine the dynamic temporal content to compensate for incomplete observations. Finally, in the composition stage, we propose to globally align the geometric context with the temporal context within a unified space by a Depth-Hypothesis-Based Transformation (DHBT) for semantic-consistent aggregation, which takes the depth hypothesis of the temporal feature volume as the distance axis and employs a voxel-pooling operation to splat the volumetric features into the unified space. This paper is structured as follows: Section II reviews related work on semantic occupancy prediction, BEV representations, and temporal modeling. Section III details our proposed Hi-SOP framework, including the hierarchical context alignment paradigm, disentangled geometric confidence-aware lifting (GCL) and temporal alignment modules, cross-frame pattern affinity (CPA) measurement, affinity-based dynamic refinement (ADR), and depth-hypothesis-based transformation (DHBT) for global composition. Section IV presents extensive experiments on SemanticKITTI, NuScenes, and NuScenes-Occupancy datasets, validating our method’s superiority in semantic scene completion and LiDAR semantic segmentation against state-of-the-art approaches. Ablation studies further analyze key components. Section V concludes the work and discusses future directions.
@@
We conduct extensive experiments to validate the advantages of our proposed hierarchical context alignment paradigm for Semantic Occupancy Prediction (SOP) in the semantic scene completion (SSC) and LiDAR semantic segmentation tasks. For SSC, our camera-based Hi-SOP outperforms state-of-the-art VoxFormer-T [li2023voxformer] on the SemanticKITTI [behley2019semantickitti] benchmark and even surpasses LiDAR-based methods on the NuScenes-Occupancy [wang2023openoccupancy] benchmark. We also evaluate our method on the NuScenes [caesar2020nuscenes] dataset for LiDAR semantic segmentation. Hi-SOP surpasses TPVFormer [huang2023tri] with a relative improvement of 24.28% in terms of mIoU. This work is an extension version that upgrades our previous ECCV-24 conference paper HTCL [li2024hierarchical2] into Hi-SOP with the following new contributions:
@@
1) Conceptually, we introduce a new hierarchical context alignment paradigm that first disentangles geometric and temporal context learning into a local-global hierarchy, and then aggregates them based on semantic consistency for a complementary composition.
@@
3) Experimentally, the initial HTCL focuses only on the semantic scene completion task, while Hi-SOP extends the framework for semantic occupancy prediction tasks including semantic scene completion and LiDAR semantic segmentation. The project website is available at https://arlo0o.github.io/hisop.github.io/.
II Related Work
II-A Semantic Occupancy Prediction
@@
Semantic Occupancy Prediction (SOP), also referred to as Semantic Scene Completion (SSC), represents a comprehensive 3D perception task that concurrently tackles semantic segmentation and scene completion [behley2019semantickitti, cai2021semantic, philion2020lift, li2023stereoscene, li2024hierarchical2]. Current studies can be broadly categorized into LiDAR-based and camera-based approaches. LiDAR-based methods have seen significant advancements driven by innovative neural architectures and intermediate representations specifically designed to address the sparsity and irregularity of LiDAR data [xia2023scpnet, zhang2018efficient, rist2021semantic, garbade2019two, cortinhal2020salsanext, zhang2020polarnet, milioto2019rangenet++]. For instance, SCPNet [xia2023scpnet] tackles challenges such as sparse input and label noise by introducing Multi-Path Blocks for multi-scale feature aggregation and a Dense-to-Sparse Knowledge Distillation strategy to enhance single-frame representation learning. SalsaNext [cortinhal2020salsanext] improves semantic segmentation of LiDAR point clouds through a residual dilated convolution stack and Bayesian uncertainty estimation, achieving state-of-the-art performance on the SemanticKITTI dataset. PolarNet [zhang2020polarnet] proposes a polar bird’s-eye-view representation to address the uneven spatial distribution of points, enabling efficient and accurate online semantic segmentation with low latency. Similarly, RangeNet++ [milioto2019rangenet++] leverages range images as an intermediate representation and introduces a novel post-processing algorithm to mitigate discretization errors, achieving real-time, high-accuracy semantic segmentation on embedded hardware. Camera-based 3D SOP has recently gained traction due to its cost-effectiveness and portability [behley2019semantickitti, silberman2012indoor, song2017semantic, straub2019replica, cai2021semantic, rist2021semantic, yan2021sparse, cheng2021s3cnet, wu2020scfusion, roldao2020lmscnet, li2020anisotropic, li2023bevdepth, cao2022monoscene, li2023voxformer, huang2023tri]. MonoScene [cao2022monoscene] pioneered the inference of geometry and semantics from a single RGB image using 2D-3D feature projection. Following this innovation, numerous studies have expanded the scope of camera-based 3D scene perception [huang2023tri, zhang2023occformer, wei2023surroundocc, li2023stereoscene]. OccFormer [zhang2023occformer] employs a monocular depth net and context net to lift voxel feature volume, which is processed with a dual-path transformer block for semantic occupancy prediction. TPVFormer [huang2023tri] introduces a tri-perspective view to enhance the detailed representation of a 3D scene. SurroundOcc [wei2023surroundocc] estimates dense 3D occupancy using multi-view image inputs. VPD [li2024time] utilizes conditional diffusion models for 3D perception tasks, including multi-view stereo and semantic occupancy prediction. In this paper, we advocate for leveraging reliable temporal data to dynamically integrate semantic context and compensate for incomplete observations.
II-B Geometry Learning in BEV Representation
@@
The bird’s-eye view (BEV) is a prevalent representation in 3D object detection, offering a comprehensive depiction of layouts and strong hallucination capabilities from a top-down perspective [philion2020lift, huang2021bevdet, li2022bevstereo, li2022bevformer, li2023bev]. The Lift-Splat approach [philion2020lift] initially introduced the extraction of BEV representations from multiple cameras by implicitly unprojecting 2D visual inputs through estimated depth distributions. To improve geometric modeling in the lifting process, BEVDepth [li2023bevdepth] employs a camera-aware monocular depth estimation module, enhancing depth accuracy in BEV-based 3D detection. As an effective representation for 3D scenarios, BEV representations have also been effectively utilized in recent occupancy-based perception works [zhang2023occformer, wei2023surroundocc, li2023stereoscene]. Notably, StereoScene [li2023stereoscene] takes advantage of stereo matching technology and utilizes stereo images to improve the geometric information in the BEV representation and achieves remarkable enhancements. In this study, we develop a new framework that incorporates monocular or stereo depth estimation to explicitly model geometric information with depth distribution confidence awareness. Moreover, the geometric context is aligned with the temporal context into a unified space through depth-hypothesis-based transformation for stable representation aggregation.
II-C Temporal Modeling in 3D Visual Perception
@@
The incorporation of temporal information has gained prominence in applications such as temporal 3D object detection [li2021enforcing, kopf2021robust, liu2023petrv2] and video depth estimation [chen2024neuralrecon, wang2023crafting], enhancing overall prediction accuracy. Temporal 3D object detection generally targets coarse-grained, regional-level predictions [liu2023petrv2, lin2022sparse4d], whereas video depth estimation techniques strive to establish correspondences across sequential video frames [long2021multi, cai2023riav]. Nevertheless, such approaches fall short in SOP, where capturing fine-grained features is crucial for dense semantic perception. VoxFormer [li2023voxformer] establishes the first temporal framework for camera-based SOP by merely stacking features from different frames, yet the temporal correspondence modeling for the dense perception task of SOP remains unexplored. GaussianWorld [zuo2025gaussianworld] proposes to exploit the scene evolution by considering the continuity of 3D occupancy. However, it fuses context from different temporal frames without explicit pattern similarity modeling, which may lead to unstable representation learning due to unaligned semantic content. GDFusion [chen2025rethinking] incorporates the formulation of RNNs to implement gradient descent for temporal information integration, considering temporal motion, temporal geometry, and scene-level temporal cues. Nevertheless, the lack of explicit temporal correspondence modeling may result in unstable learning, and the RNN-based formulation is susceptible to cumulative errors [li2017diffusion, mao2022review, li2024time]. In contrast, our approach explicitly models temporal context correlation through homography warping and affinity-aware dynamic refinement, enabling reliable aggregation of aligned temporal content while mitigating the impact of incomplete observations.
@@
III Methodology
III-A Overview
III-A1 Preliminary
Given a sequence of temporal RGB images $I_{{set}}^{{rgb}}=\{I_{t}^{{rgb}},I_{t-1}^{{rgb}},\ldots\}$, our objective is to estimate the semantic voxel grid for semantic scene completion or LiDAR segmentation [cao2022monoscene, zhang2023occformer]. We focus on current and historical image frames, excluding future frames [li2023voxformer] to devise a practical method for real-world applications. The scene is represented as a voxel grid $\mathbf{V}$ with dimensions $\mathbb{R}^{H\times W\times Z}$, where $H$, $W$, and $Z$ denote the height, width, and depth of the grid, respectively. Each voxel within this grid is classified into one of the semantic categories in the set $\{c_{0},c_{1},\ldots,c_{N}\}$, where $c_{0}$ indicates empty space and $\{c_{1},c_{2},\ldots,c_{N}\}$ correspond to $N$ distinct semantic classes. With the proposed framework $\Theta$, we seek to learn a transformation defined as: